# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ V-C1 Proof of Lemma 7
$\|{\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)\|_{\infty}\leq c\sqrt{\frac{d_{\max}\log n}{k}}\leq c\frac{d_{\max}}{\sqrt{k}}.$ (74)
Substituting the above bound in Eq. 73 we obtain the bound in Eq. 72 for some constant $\tilde{c}_{2}$.
Now again we need to show that $\|\hat{P}-P\|_{2}\leq O(\sqrt{d_{\max}/k})$ with high probability. In case of complete graph, by [80, Theorem 4.4.5], we have $\|\hat{P}-P\|_{2}\leq 6c_{p}\sqrt{d_{\max}/k}$ with high probability for some constant $c_{p}$ by the same argument as (69) (since ${\mathrm{var}}(\hat{p}_{ij})\leq 1/k$). For a general graph model, and by application of matrix Bernstein inequality, again a $\log n$ factor becomes a bottleneck as we obtain $\|\hat{P}-P\|_{2}\leq O\big{(}\sqrt{\frac{d_{\max}\log n}{k}}\big{)}$ (with high probability). Therefore, we again utilize [81, Corollary 2.15] to obtain tighter concentration bounds on $\|\hat{P}-P\|_{2}$. Applying, Lemma 14 we obtain
$\mathbb{P}\left(\|\hat{P}-P\|_{2}\geq\sqrt{\frac{d_{\max}}{k}}+c\left(\frac{d_{\max}^{1/4}}{\sqrt{4k}}(\log n)^{3/4}+\sqrt{\frac{t}{4k}}+\frac{d_{\max}{}^{1/3}}{(2k)^{2/3}}t^{2/3}+\frac{t}{k}\right)\right)\leq 4ne^{-t}.$ (75)
Now combining Eqs. 65, 70, 71 and 75, we obtain
$\mathbb{P}\left(T_{1}\geq c_{1}\frac{n\|\pi\|_{\infty}^{2}}{k}+\frac{\tilde{c}n\|\pi\|_{\infty}^{2}}{k}\left(\left(\frac{(\log n)^{3}}{d_{\max}}\right)^{1/4}+\sqrt{\frac{t}{d_{\max}}}+\left(\frac{t^{4}}{kd_{\max}}\right)^{1/6}+\sqrt{\frac{t^{2}}{kd_{\max}}}\right)\right)\leq 8ne^{-t}+O\bigg{(}\frac{1}{n^{3}}\bigg{)},$ (76)
where the term $O(\frac{1}{n^{3}})$ is added to account for the probability with which the event $\mathcal{A}_{2}$ and Eqs. 68 and 74 holds. Substituting $t=\bar{c}\log n$ for some constant $\bar{c}$, and utilizing the fact that $d_{\max}\geq(\log n)^{4}$, we obtain that there exists a constant $c_{\alpha}$ such that with probability at least $1-O(1/n^{3})$, we have $T_{1}\leq c_{\alpha}\frac{n\|\pi\|_{\infty}^{2}}{k}$. The corresponding lower bounds on $T_{1}$ can be obtained in a similar manner, and therefore, for the tail bound on $|T_{1}|$, we get a factor of $2$, thus proving Eq. 56.
@@
$T_{2}$ $=\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)^{2}-\left(\pi_{i}+\pi_{j}\right)^{2}\right)p_{ij}^{2}+\hat{\pi}_{j}^{2}-\pi_{j}^{2}-2\left(\hat{\pi}_{j}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)-\pi_{j}\left(\pi_{i}+\pi_{j}\right)\right)p_{ij}$
$=\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\hat{\pi}_{i}^{2}-\pi_{i}^{2}\right)p_{ij}^{2}+\left(\hat{\pi}_{j}^{2}-\pi_{j}^{2}\right)\left(p_{ij}^{2}+1-2p_{ij}\right)+2\left(\hat{\pi}_{i}{\hat{\pi}}_{j}-\pi_{i}\pi_{j}\right)\left(p_{ij}^{2}-p_{ij}\right)$
$=\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\hat{\pi}_{i}-\pi_{i}\right)\left(\hat{\pi}_{i}+\pi_{i}\right)\left(p_{ij}^{2}+\left(1-p_{ji}\right)^{2}\right)+2\left({\hat{\pi}}_{i}-\pi_{i}\right)\left(\hat{\pi}_{j}-\pi_{j}\right)\left(p_{ij}^{2}-p_{ij}\right)$
@@
$\left({\pi}_{i}(p_{ij}^{2}+\left(1-p_{ji}\right)^{2})+\pi_{j}\left(p_{ji}^{2}-p_{ji}\right)+\pi_{j}\left(p_{ij}^{2}-p_{ij}\right)\right)+2\left({\hat{\pi}}_{i}-\pi_{i}\right)\left(\hat{\pi}_{j}-\pi_{j}\right)\left(p_{ij}^{2}-p_{ij}\right)$
@@
$\ \ \sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}2\left(\hat{\pi}_{i}-\pi_{i}\right)p_{ij}\left({\pi}_{i}p_{ij}+\pi_{j}p_{ij}-\pi_{j}\right)+2\left(\hat{\pi}_{i}-\pi_{i}\right)\left(1-p_{ji}\right)\left(\pi_{i}-\pi_{i}p_{ji}-\pi_{j}p_{ji}\right)$
@@
$\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}\|\hat{\pi}-\pi\|^{2}_{2}\|P_{3}{\mathbf{1}}_{n}\|_{\infty}+4\sqrt{d_{\max}}\|{\hat{\pi}}-\pi\|_{2}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}+\|{\hat{\pi}}-\pi\|^{2}_{2}\lambda_{1}(P-P_{2})$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\leq}}\frac{4c_{2}\|\pi\|_{2}}{\sqrt{k}}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}+2\frac{c^{2}_{2}\|\pi\|^{2}_{2}}{k}$
$\stackrel{{\scriptstyle}}{{\leq}}\frac{4c_{2}\sqrt{n}\|\pi\|_{\infty}}{\sqrt{k}}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}+\frac{2nc^{2}_{2}\|\pi\|^{2}_{\infty}}{k},$
where the explanation for the middle term of $\zeta_{1}$ is provided below and $\zeta_{2}$ follows from since the event $\mathcal{A}_{2}$ holds and the fact that $P-P_{2}$ is a non-negative matrix; therefore, all of its eigenvalues are smaller than its spectral radius in absolute value. Therefore, by Gershgorin circle theorem, $\lambda_{1}(P-P_{2})\leq d_{\max}$. The middle term in $\zeta_{1}$ is bounded as
$\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}2\left(\hat{\pi}_{i}-\pi_{i}\right)p_{ij}\left({\pi}_{i}p_{ij}+\pi_{j}p_{ij}-\pi_{j}\right)$ $=\sum_{i=1}^{n}2\left(\hat{\pi}_{i}-\pi_{i}\right)\sum_{\begin{subarray}{c}j:(i,j)\in{\mathcal{E}},\\ j\neq i\end{subarray}}p_{ij}\left({\pi}_{i}p_{ij}+\pi_{j}p_{ij}-\pi_{j}\right)$
$\leq\|\hat{\pi}-\pi\|_{2}\|(\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})){\mathbf{1}}_{n}\|_{2}$
$\leq\sqrt{d_{\max}}\|\hat{\pi}-\pi\|_{2}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}},$
where last inequality follows by utilizing the fact at most $d_{\max}$ entries are non-zero in any row of the matrix $\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}}),$ and the fact that $(\sum_{i=1}^{n}a_{i})^{2}\leq(\sum_{i=1}^{n}|a_{i}|)^{2}\leq n\sum_{i=1}^{n}a_{i}^{2}$. This holds because the absolute value of each entry of $P_{2}-P$ and $P_{3}$ is upper bounded by 1. Similarly, we can obtain the corresponding lower bounds on $T_{2}$ in the same fashion as above, and this completes the proof.
@@ V-C2 Proof of Lemma 8
${\mathbb{E}}[T_{3}]$ $=\sum_{i=1}^{n}\sum_{j=1}^{n}(\pi_{i}+\pi_{j})^{2}\frac{{\mathbb{E}}[Z_{ij}^{2}]-{\mathbb{E}}[Z_{ij}]}{k_{ij}(k_{ij}-1)}+\pi_{j}^{2}-2(\pi_{i}+\pi_{j})\pi_{j}\frac{{\mathbb{E}}[Z_{ij}]}{k_{ij}}$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}(\pi_{i}+\pi_{j})^{2}p_{ij}^{2}+\pi_{j}^{2}-2(\pi_{i}+\pi_{j})\pi_{j}p_{ij}=\|\Pi P-P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|^{2}_{\mathrm{F}}.$
@@
${\mathrm{var}}\left(T_{3}\right)$ $=\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\pi_{i}+\pi_{j}\right)^{4}{\mathrm{var}}(\hat{Y}_{ij})+4\pi_{j}^{2}\left(\pi_{i}+\pi_{j}\right)^{2}{\mathrm{var}}(\hat{p}_{ij})-4\left(\pi_{i}+\pi_{j}\right)^{3}\pi_{j}\left({\mathbb{E}}[\hat{Y}_{ij}\hat{p}_{ij}]-{\mathbb{E}}[\hat{Y}_{ij}]{\mathbb{E}}[\hat{p}_{ij}]\right)$
=$\stackrel{{\scriptstyle\zeta_{1}}}{{=}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\pi_{i}+\pi_{j}\right)^{4}\left(\frac{2p_{ij}^{2}+4(k-2)p_{ij}^{3}+(6-4k)p_{ij}^{4}}{k(k-1)}\right)+\frac{4\pi_{j}^{2}\left(\pi_{i}+\pi_{j}\right)^{2}p_{ij}\left(1-p_{ij}\right)}{k}$
@@
$\stackrel{{\scriptstyle}}{{=}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\frac{2(\pi_{i}+\pi_{j})^{2}}{k}\bigg{(}\frac{(\pi_{i}+\pi_{j})^{2}(p_{ij}^{2}-2p_{ij}^{3}+p_{ij}^{4})}{k-1}+2(\pi_{i}+\pi_{j})^{2}p_{ij}^{3}(1-p_{ij})$
@@
$\stackrel{{\scriptstyle\zeta_{2}}}{{\leq}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\frac{2(\pi_{i}+\pi_{j})^{2}}{k}\bigg{(}\frac{(\pi_{i}+\pi_{j})^{2}p_{ij}^{2}(1-p_{ij})^{2}}{k-1}+\frac{1}{2}((\pi_{i}+\pi_{j})p_{ij}-\pi_{j})^{2}\bigg{)}$
$\stackrel{{\scriptstyle\zeta_{3}}}{{\leq}}\frac{4\|\pi\|_{\infty}^{2}}{k}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}^{2}+\frac{32\|\pi\|_{\infty}^{4}nd_{\max}}{k(k-1)}\times\frac{1}{16}$
@@
This completes the proof.

### v2

@@ V-C1 Proof of Lemma 7
$\|{\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)\|_{\infty}\leq c\sqrt{\frac{d_{\max}\log n}{k}}.$ (72)
Substituting the above bound in Eq. 71 and using the same argument as in Eq. 66, we obtain the bound in Eq. 70 for some constant $\tilde{c}$.
Now again we need to show that $\|\hat{P}-P\|_{2}\leq O(\sqrt{d_{\max}/k})$ with high probability. In case of complete graph, by [HDP, Theorem 4.4.5], we have $\|\hat{P}-P\|_{2}\leq 6c_{p}\sqrt{d_{\max}/k}$ with high probability for some constant $c_{p}$ by the same argument as (67) (since ${\mathrm{var}}(\hat{p}_{ij})\leq 1/k$). For a general graph model, and by application of matrix Bernstein inequality, again a $\log n$ factor becomes a bottleneck as we obtain $\|\hat{P}-P\|_{2}\leq O\big(\sqrt{\frac{d_{\max}\log n}{k}}\big)$ (with high probability). Therefore, we again utilize [brailovskaya2022, Corollary 2.15] to obtain tighter concentration bounds on $\|\hat{P}-P\|_{2}$. Applying, Lemma 14 we obtain
[…]
Now combining Eqs. 62, 68, 69 and 73, we obtain
[…]
$\left.\left.\!\sqrt{\frac{t}{d_{\max}}}+\left(\frac{t^{4}}{kd_{\max}}\right)^{\frac{1}{6}}\!\!+\sqrt{\frac{t^{2}}{kd_{\max}}}\right)\right)\leq 8ne^{-t}+O\bigg(\frac{1}{n^{3}}\bigg),$ (74)
where the term $O\big(\frac{1}{n^{3}}\big)$ is added to account for the probability with which the event $\mathcal{A}_{2}$, Eq. 65, and Eq. 72 hold. Substituting $t=\bar{c}\log n$ for some constant $\bar{c}$, and utilizing the fact that $d_{\max}\geq(\log n)^{4}$, we obtain that there exists a constant $c_{\alpha}$ such that with probability at least $1-O(1/n^{3})$, we have $T_{1}\leq c_{\alpha}\frac{n\|\pi\|_{\infty}^{2}\sqrt{h}}{k\xi^{2}}$. The corresponding lower bounds on $T_{1}$ can be obtained in a similar manner, and therefore, for the tail bound on $|T_{1}|$, we get a factor of $2$ on the tail-probability, thus proving Eq. 54.
@@
$T_{2}$ $=\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)^{2}-\left(\pi_{i}+\pi_{j}\right)^{2}\right)p_{ij}^{2}+\hat{\pi}_{j}^{2}-\pi_{j}^{2}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ -2\left(\hat{\pi}_{j}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)-\pi_{j}\left(\pi_{i}+\pi_{j}\right)\right)p_{ij}$
$=\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\hat{\pi}_{i}^{2}-\pi_{i}^{2}\right)p_{ij}^{2}+\left(\hat{\pi}_{j}^{2}-\pi_{j}^{2}\right)\left(p_{ij}^{2}+1-2p_{ij}\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ +2\left(\hat{\pi}_{i}{\hat{\pi}}_{j}-\pi_{i}\pi_{j}\right)\left(p_{ij}^{2}-p_{ij}\right)$
$=\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\hat{\pi}_{i}-\pi_{i}\right)\left(\hat{\pi}_{i}+\pi_{i}\right)\left(p_{ij}^{2}+\left(1-p_{ji}\right)^{2}\right)$
$\ \ \ \ \ \ \ \ \ \ \ +2\left({\hat{\pi}}_{i}-\pi_{i}\right)\left(\hat{\pi}_{j}-\pi_{j}\right)\left(p_{ij}^{2}-p_{ij}\right)$
@@
$\left({\pi}_{i}(p_{ij}^{2}+\left(1-p_{ji}\right)^{2})+\pi_{j}\left(p_{ji}^{2}-p_{ji}\right)+\pi_{j}\left(p_{ij}^{2}-p_{ij}\right)\right)$
$\ \ \ \ \ \ \ +2\left({\hat{\pi}}_{i}-\pi_{i}\right)\left(\hat{\pi}_{j}-\pi_{j}\right)\left(p_{ij}^{2}-p_{ij}\right)$
@@
$\ \ \ \ \ \ \ \ \ \ \ \sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}2\left(\hat{\pi}_{i}-\pi_{i}\right)p_{ij}\left({\pi}_{i}p_{ij}+\pi_{j}p_{ij}-\pi_{j}\right)$
$\ \ \ \ \ \ \ \ \ \ +2\left(\hat{\pi}_{i}-\pi_{i}\right)\left(1-p_{ji}\right)\left(\pi_{i}-\pi_{i}p_{ji}-\pi_{j}p_{ji}\right)$
@@
$\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}\|\hat{\pi}-\pi\|^{2}_{2}\|P_{3}{\mathbf{1}}_{n}\|_{\infty}$
$\ \ \ \ \ \ \ \ \ +4\sqrt{d_{\max}}\|{\hat{\pi}}-\pi\|_{2}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}$
$\ \ \ \ \ \ \ \ \ \ \ +\|{\hat{\pi}}-\pi\|^{2}_{2}\|P-P_{2}\|_{2}$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\leq}}\frac{4c_{2}\sqrt{n}\|\pi\|_{\infty}}{\sqrt{k}}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}\!+\!\frac{3c^{2}_{2}n\|\pi\|^{2}_{\infty}}{k},$
where the explanation for the middle term of $\zeta_{1}$ is provided below, and $\zeta_{2}$ follows since the event $\mathcal{A}_{2}$ holds and $P-P_{2}$ is a non-negative matrix, which implies that
$\|P-P_{2}\|_{2}$ $\leq\sqrt{\max_{i\in[n]}\|e_{i}^{\mathrm{T}}(P-P_{2})\|_{1}}\sqrt{\max_{j\in[n]}\|(P-P_{2})e_{j}\|_{1}}$
$\leq d_{\max}.$
Moreover, we have $\|P_{3}{\mathbf{1}}_{n}\|_{\infty}\leq 2d_{\max}$. The middle term in $\zeta_{1}$ is bounded as
$\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}2\left(\hat{\pi}_{i}-\pi_{i}\right)p_{ij}\left({\pi}_{i}p_{ij}+\pi_{j}p_{ij}-\pi_{j}\right)$
$\ \ \ \ \ \ \ =\sum_{i=1}^{n}2\left(\hat{\pi}_{i}-\pi_{i}\right)\sum_{\begin{subarray}{c}j:(i,j)\in{\mathcal{E}},\\ j\neq i\end{subarray}}p_{ij}\left({\pi}_{i}p_{ij}+\pi_{j}p_{ij}-\pi_{j}\right)$
$\ \ \ \ \ \ \ \leq 2\sqrt{d_{\max}}\|\hat{\pi}-\pi\|_{2}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}},$
where last inequality follows by utilizing the fact at most $d_{\max}$ entries are non-zero in any row of the matrix $\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}}),$ and the fact that $(\sum_{i=1}^{n}a_{i})^{2}\leq(\sum_{i=1}^{n}|a_{i}|)^{2}\leq n\sum_{i=1}^{n}a_{i}^{2}$. This holds because the absolute value of each entry of $P_{2}-P$ and $P_{3}$ is upper bounded by 1. Similarly, we can obtain the corresponding lower bounds on $T_{2}$ in the same fashion as above, and this completes the proof. ∎
@@ V-C2 Proof of Lemma 8
${\mathbb{E}}[T_{3}]$ $=\sum_{i=1}^{n}\sum_{j=1}^{n}(\pi_{i}+\pi_{j})^{2}\frac{{\mathbb{E}}[Z_{ij}^{2}]-{\mathbb{E}}[Z_{ij}]}{k_{ij}(k_{ij}-1)}+\pi_{j}^{2}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ -2(\pi_{i}+\pi_{j})\pi_{j}\frac{{\mathbb{E}}[Z_{ij}]}{k_{ij}}$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}(\pi_{i}+\pi_{j})^{2}p_{ij}^{2}+\pi_{j}^{2}-2(\pi_{i}+\pi_{j})\pi_{j}p_{ij}$
$=\|\Pi P-P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|^{2}_{\mathrm{F}}.$
@@
${\mathrm{var}}\left(T_{3}\right)\!=\!\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\!\left(\pi_{i}+\pi_{j}\right)^{4}{\mathrm{var}}(\hat{Y}_{ij})+4\pi_{j}^{2}\left(\pi_{i}+\pi_{j}\right)^{2}{\mathrm{var}}(\hat{p}_{ij})$
$\ \ \ \ \ \ \ \ -4\left(\pi_{i}+\pi_{j}\right)^{3}\pi_{j}\left({\mathbb{E}}[\hat{Y}_{ij}\hat{p}_{ij}]-{\mathbb{E}}[\hat{Y}_{ij}]{\mathbb{E}}[\hat{p}_{ij}]\right)$
$\stackrel{{\scriptstyle\zeta_{1}}}{{=}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\pi_{i}+\pi_{j}\right)^{4}\left(\frac{2p_{ij}^{2}+4(k-2)p_{ij}^{3}+(6-4k)p_{ij}^{4}}{k(k-1)}\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{4\pi_{j}^{2}\left(\pi_{i}+\pi_{j}\right)^{2}p_{ij}\left(1-p_{ij}\right)}{k}$
@@
$\stackrel{{\scriptstyle}}{{=}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\frac{2(\pi_{i}+\pi_{j})^{2}}{k}\bigg(\frac{(\pi_{i}+\pi_{j})^{2}(p_{ij}^{2}-2p_{ij}^{3}+p_{ij}^{4})}{k-1}$
$\ \ \ \ \ \ \ \ \ \ \ +2(\pi_{i}+\pi_{j})^{2}p_{ij}^{3}(1-p_{ij})$
@@
$\stackrel{{\scriptstyle\zeta_{2}}}{{\leq}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\frac{2(\pi_{i}+\pi_{j})^{2}}{k}\bigg(\frac{(\pi_{i}+\pi_{j})^{2}p_{ij}^{2}(1-p_{ij})^{2}}{k-1}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{2}((\pi_{i}+\pi_{j})p_{ij}-\pi_{j})^{2}\bigg)$
$\stackrel{{\scriptstyle\zeta_{3}}}{{\leq}}\frac{4\|\pi\|_{\infty}^{2}}{k}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}^{2}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{32\|\pi\|_{\infty}^{4}nd_{\max}}{k(k-1)}\times\frac{1}{16}$
@@
This completes the proof. ∎